OpenStudy (anonymous):

Write in slope-intercept form: 2x + 5y= 15

OpenStudy (anonymous):

slope-intercept form: y = mx + b where m is the slope and b in the y-intercept. simply re-arrange them; first subtract 2x from both sides: 5y = 15 - 2x divide both sides by 5 y = 3 - (2x/5)

OpenStudy (anonymous):

thank you! :)